OpenStudy (anonymous):

How do I determine the area of a region bounded by a function and the x-axis if given y=cosx and interval is [0,π/2]? This has to do with the Fundamental Theorem of Calculus. Thanks in advance :)

OpenStudy (amistre64):

you integrate the function from 0 to pi/2

OpenStudy (amistre64):

\[\int_{0}^{pi/2}f(x)dx=\left.F(x)\right|_{0}^{pi/2}\] \[F(pi/2)-F(0)=Area.under.curve\]

OpenStudy (amistre64):

do you know where cos(x) is derived from?

OpenStudy (amistre64):

F(x) = ? F'(x) = f(x) = cos(x)

OpenStudy (anonymous):

I don't know what you mean. cos(x) is given. I tried following the steps on the attached file but didn't understand how to get the "sin(x) on the second step. (please refer to attachment) Thanks.

OpenStudy (amistre64):

there are 2 terms that are used for this, integration is one of them; and anti-differentation is the other. Which one sounds more familiar to you?

OpenStudy (anonymous):

Well they are the same thing, right?

OpenStudy (amistre64):

yes, but since your befuddled about what ive said i have to make sure i relay it to you in terms that you are capable of associating with

OpenStudy (amistre64):

y = cos(x) might as well be read as: y' = cos(x), find the y that got dervied

OpenStudy (amistre64):

derived

OpenStudy (amistre64):

in other words, underive cos(x)

OpenStudy (anonymous):

so y=cosx is not f(x)? It is f'(x)?

OpenStudy (amistre64):

that is the best way to view it yes. Call it by a name that you can see it for what it truly is.

OpenStudy (amistre64):

f'(x) = cos(x) is what you want to undo to get back to f(x)

OpenStudy (anonymous):

but then wouldn't f(x) be -sinx, not sinx, because the derivative is y' cos(x)=-sin(x)?

OpenStudy (amistre64):

youre going the wrong way ..... let me try to draw it out :